# Taylor’s hypothesis and its impact on flux measurements in a forest clearcut

### v1

@@ 3.3.2 A critical frequency limit
Despite the limitations of DTS measurements, we test one hypothesis of Cheng et al. (2017) where they proposed that by using scale-dependent convection speeds one can increase the flux contributions from smaller time scales. This eventually could serve as a correction for improving the imbalance in surface energy components as commonly being observed in micrometeorological measurements (Mauder et al., 2020). In contrast to ours, Cheng et al. (2017) found that the scale-dependent convection speeds of temperature structures vary as ${(\kappa z)}^{-1/3}$. If we use that information to convert the frequencies to wavenumbers, then after some algebraic manipulation, one can show that the wavenumbers ($\kappa z$) can be estimated from frequencies ($f$) by solving the polynomial,
${(\kappa z)}^{3}=a(1+\kappa z),$ (13)
where $a=(2\pi fz/U_{C})^{3}$. To derive Eq. 13, we have assumed that the convection speeds can be expressed as,
$\frac{U_{c}(\kappa z)}{U_{c}}=\frac{1}{{(1+\kappa z)}^{1/3}},$ (14)
which appears to be supported by the plot shown in Fig. S4b of the Supplementary Material.
In Figs. 8c-d, we show the premultiplied heat and moisture cospectra ($fS_{wx}(f)$, where $x=T,q$) where the frequencies are converted to wavenumbers using mean wind speed ($U$, red line), bulk convective speed ($U_{C}$, magenta line), and scale-dependent convective speeds ($U_{C}(\kappa z)$) from our observations (pink line) and from Cheng et al. (2017) (blue line). We adopt the premultiplied form since the area under this log-linear coordinate system is directly related to the flux contributions across different scales (Kaimal and Finnigan, 1994). The results from Figs. 8c-d clearly demonstrate that using our scale-dependent convective speeds (Eq. 9) cause an appreciable increase in the heat and moisture fluxes at smaller scales of the flow. More importantly, the differences between the cospectral shapes when using either $U$, $U_{C}$, or $U_{C}(\kappa z)$ relationship from Cheng et al. (2017) to convert the frequencies to wavenumbers are small.
@@
The transformations between wavenumbers and frequencies (Eqs. 11 and 13) also yield an unrealistic outcome since the wavenumbers estimated from both Eqs. 11 and 13 are strong non-linear functions of frequency. If these functions are plotted for a specific 30-min period, as done in Fig. 8e, it is apparent that the wavenumbers estimated from Eq. 11 can be up to $10^{6}\rm\ m^{-1}$ when converting the EC sampling frequency (10 Hz) to wavenumber. This cannot be true as it suggests that the EC systems can measure eddies comparable to the dissipation scales, which are clearly smaller than the path length (distance between the transducers) of the sonic anemometers themselves. For our sonic anemometer system, the path length is around 16 cm, and hence the wavenumbers that can be realistically resolved is around $2\pi/0.3\rm\ m^{-1}$. To avoid the aliasing effect, the minimum wavelength being resolved is assumed to be twice the path length (i.e. 30 cm). This limit is shown as a horizontal dash-dotted line in Fig. 8e. If we call this minimum wavelength being set by the physical dimension of the sonic anemometer as $\kappa_{\rm cr}$, then from Eq. 11, one can convert $\kappa_{\rm cr}$ to a frequency limit ($f_{\rm cr}$) in the temporal space. Accordingly, one can estimate $f_{\rm cr}$ as,
$f_{\rm cr}=\frac{\kappa_{\rm cr}U_{C}}{2\pi{(1+\kappa_{\rm cr}z)}^{4/5}}.$ (15)
@@
For our set up, $z=3.1$ m, $\kappa_{\rm cr}=2\pi/0.3\rm\ m^{-1}$, and $U_{C}$ can be found from the space-time correlation curves (see Fig. 4). For the along-wind cases, $U_{C}$ values are averaged over all the five sections. By replacing all these values in Eq. 15, $f_{\rm cr}$ appears to be $\approx$ 0.3 Hz (for both along- and perpendicular-wind periods), with some run-to-run variations since $U_{C}$ values are not fixed.
3.3.3 High-frequency flux uncertainties
If we consider the $U_{C}(\kappa z)$ values and therefore the derived wavenumbers from Eq. 11 to be correct, then it is unclear what the EC observations represent at frequencies above $f_{\rm cr}$ (i.e. $f\geq f_{\rm cr}$). This is because, beyond $f_{\rm cr}$, the eddy length scales are smaller than the physical dimension of the instrument itself. Of course this problem is not that severe if the wavenumbers are estimated from Eq. 13, i.e. using the Cheng et al. (2017) relationship (see the blue line in Fig. 8e). By following the same procedure, $f_{\rm cr}$ values for the Cheng et al. (2017) relationship can be derived as,
$f_{\rm cr}=\frac{\kappa_{\rm cr}U_{C}}{2\pi{(1+\kappa_{\rm cr}z)}^{1/3}}.$ (16)
If one substitutes all the values from our set up in Eq. 16, $f_{\rm cr}$ appears to be $\approx$ 2 Hz. In general, $f_{\rm cr}$ values are dependent on the power-law exponent ($\alpha$) of the scale-dependent convection speeds, $U_{C}(\kappa z)/U_{C}=1/(1+\kappa z)^{\alpha}$, with $10<\alpha<$ 1. Here, $\alpha=1$ corresponds to a limit for the randomly-shuffled samples while $\alpha=0$ indicates no scale-dependency.
The existence of $f_{\rm cr}$ ultimately results in a situation where the fluxes associated with EC systems at higher frequencies are largely uncertain. In a fractional form, this high-frequency flux uncertainty can be expressed as,
[…]
where $S_{wx}$ is the cospectral density between $w$ and $x$. Here $x$ can be any constituent, e.g. $u$, $T$, H2O or CO2. For any EC system, this fraction can be estimated by computing $f_{\rm cr}$ from either Eq. 15 or Eq. 16, depending on the flow situations. In the absence of any concrete evidence, we hypothesize that Eq. 15 is valid for highly-heterogeneous flows (e.g., forest clearcut) while Eq. 16 can be used for homogeneous surface layer flows. For highly-heterogeneous conditions, $U_{C}$ can be obtained from Eq. 7.
For our case, the fraction of fluxes within the frequency range $0<f<f_{\rm cr}$ is around 85-90% of the total flux being resolved. Thus, between 10-15% of the flux is related to the frequency range that is questionable given the physical dimension of the sonic-anemometer (see above). This has significant impact on the long-term estimations of ecosystem gross primary productivity, net ecosystem exchange, evapotranspiration, sensible heat and momentum fluxes from the EC systems. Potentially, Eq. 17 could be extended to reevaluate the estimations of velocity and scalar variances and turbulence kinetic energy, and therefore has implications for surface layer parameterizations. One other impact of Eq. 17 is on the surface energy balance.
[…]
@@ 3.3.3 High-frequency flux uncertainties
From a practical standpoint, since all our results are from a single height, we cannot establish whether $z$ is the ideal scaling variable to describe $U_{c}(\kappa)$. If $z$ indeed is the correct scaling parameter, then depending on the observation height, $f_{\rm cr}$ values will change (see Eqs. 15 and 16). Accordingly, the high-frequency flux uncertainty associated with the EC systems would also change, and possibly will increase at higher heights since $f_{\rm cr}$ is inversely proportional to $z$. We present our conclusions in the next section.
@@ 4 Conclusion
In this study, we have developed a comprehensive framework to test the validity of Taylor’s hypothesis and its impact on turbulence statistics over a highly heterogeneous forest clearcut surface. To execute this objective, an extensive dataset encompassing both Distributed Temperature Sensing (DTS) and Eddy Covariance (EC) measurements were used. These datasets were collected during a five-month field campaign (May-September 2024) and were separated into two different categories. First, when the mean wind speed was parallel to the forest edge, and second, when it was perpendicular to the edge. For the former, our results demonstrated that the Taylor’s hypothesis remained invalid as the turbulent intensities of the streamwise wind fluctuations exceeded 0.5. An empirical relationship is proposed to estimate the convective speeds in such extreme conditions from the sonic anemometer observations. Corresponding to the latter scenario, no such relationship was apparent since depending on whether the wind approached from the forest or from the clearcut the results were different.
Despite this discrepancy, for both along- and perpendicular-wind cases, the ratios of the convective speeds of positive and negative temperature structures remained at a constant value of 0.8, irrespective of the temperature skewness measured by the EC system. We used this information to compute the sizes of up- and down-plumes in a convective boundary layer. We found that the sizes of the down-plumes can be as large as 12 times the measurement height and therefore would require considerably larger area for their statistics to converge. This result informs about what should be the appropriate horizontal domain size while setting up the large eddy simulations of convective boundary layers. Furthermore, the different convection speeds for the positive and negative temperature structures suggests that an EC tower sampling the flow over time produces a biased sample of the flow when compared to sampling taken over space.
We demonstrated that the convective speeds of temperature structures at inertial subrange scales behave in a power-law fashion with increasing wavenumbers. These wavenumbers were directly computed from the DTS spatial data. Contrary to a previous study over a homogeneous atmospheric surface layer flow (Cheng et al., 2017), the exponent of this power law was $-0.8$ instead of $-1/3$. The $-1/3$ exponent was theoretically derived under the assumption of homogeneous and isotropic turbulence at smaller scales of the flow. For the present case in hand, we hypothesize that the heterogeneity in the surface conditions violate these assumptions even at smaller scales of the flow. We evaluated the impact of this revised scale-dependent convection speed on the turbulence spectra of velocity and scalars (temperature, moisture, and $\rm CO_{2}$), estimated from the temporal EC measurements. The results displayed an extensive $-5/3$ spectral slope (over two decades of wavenumbers) for the streamwise velocity component and the three different scalars when the frequencies were converted to wavenumbers using our scale-dependent convective speeds. For the flux cospectra, instead of a $-7/3$ slope, a $-5/3$ slope was recovered under such conversion.
@@
The appearance of a slope flatter than $-7/3$ indicates an enhancement of turbulent fluxes at smaller scales of the flow. However, the non-linear conversion of frequencies to wavenumbers through scale-dependent convective speeds generates unrealistically large wavenumber values. The physical dimension of the sonic anemometers sets a strict constraint on the wavenumbers that can be resolved by the EC systems. In a temporal domain, this eventually results in a critical frequency limit beyond which the EC measurements cannot be trusted. This critical limit depends on the measurement height and on the power-law exponent of the scale-dependent convective speeds. The existence of this limit dictates that there is a high-frequency flux uncertainty associated with EC systems, which, in turn, has an adverse effect on the estimation of surface energy budget. Before we outline our future research directions, Hilland and Christen (2024) asked whether the Taylor’s hypothesis is ever satisfied in atmospheric flows. For a flat desert surface, Hilland and Christen (2024) found that the Taylor’s hypothesis appeared to be valid only between spatial scales 5 to 16 m. Over a forest clearcut, our results suggest that for the periods when the wind approached the clearcut location from the forest, the Taylor’s hypothesis remained valid at scales $\kappa z<1$, where $\kappa$ is the streamwise wavenumber and $z$ is the measurement height. This was realized notwithstanding the fact that during such periods the flow was disturbed as it encountered abrupt changes in the surface roughness.
@@
At present, these results are representative of buoyancy-driven flows in a heterogeneous location and are also limited by a single observation height. One another drawback is, the estimations of the convective speeds are based on an imperfect temperature signal and we inherently assume that the same could be extended to other scalars and velocity components. It is not immediately apparent whether the structures in all these different turbulent variables move at same speeds. Therefore, for future studies, more extensive spatial data are needed to verify our framework, which ideally should span multiple heights, stability conditions (stable, neutral, and convective), surface roughness (forest canopies, grasslands, arid deserts, water-bodies, etc.), and signal types (velocity and scalars). To achieve this ambitious goal, the observations alone will not suffice and large eddy simulations would be needed to complement the DTS measurements.
@@ Appendix A Assessment of stability
In this appendix, we report the statistics of velocity and temperature fluctuations from the EC measurements corresponding to the along-wind conditions. The conclusions remain the same for perpendicular-wind conditions and hence not being shown. These statistics are used to assess the stability conditions. In practice, the Obukhov similarity parameter is commonly used to determine the atmospheric stability, but we deliberately refrained from that approach since this theory was developed for homogeneous flows and its validity is questionable over heterogeneous surfaces such as a forest clearcut. In Fig. 9a, we show the PDFs of velocity and temperature fluctuations. Typical to a buoyancy-driven flow, the PDFs of temperature fluctuations remain strongly non-Gaussian as opposed to the velocity ones ($u^{\prime}$ and $w^{\prime}$). In fact, the PDFs of temperature fluctuations can be represented through a cumulant expansion method, otherwise known as Gram-Charlier distribution (Chowdhuri and Prabha, 2019). This distribution can be expressed as:
@@
where $\mathcal{S}(T)$ is the skewness of temperature fluctuations. The non-Gaussianity in temperature statistics is also reflected in how the JPDFs between the vertical velocity and temperature fluctuations behave. From Fig. 9b, it is apparent that the JPDFs are strongly skewed towards quadrant III, which represents the downdraft motions. To further investigate the heat transport characteristics for our selected runs, we computed the flux-weighted JPDFs, expressed as, $w^{\prime}T^{\prime}P(w^{\prime},T^{\prime})$, where $P(w^{\prime},T^{\prime})$ is the JPDF between $w^{\prime}$ and $T^{\prime}$. The contours of this quantity are plotted in Fig. 9c, and the area under them is proportional to the amount of heat being transported (Nakagawa and Nezu, 1977). It is immediately clear from Fig. 9c that the transport of heat is strongly dominated by the down-gradient quadrants (I and III) with very small amounts being transported in counter-gradient directions (II and IV).
This organization results in a situation where the heat is transported more efficiently than momentum. One could verify this argument in Fig. 9d, where the correlation coefficients between $w^{\prime}$ and $T^{\prime}$ ($R_{wT}$) remain considerably larger than the ones between $w^{\prime}$ and $u^{\prime}$ ($R_{wu}$). Since $R_{wu}$ values are negative, their absolutes ($|R_{wu}|$) are used on the $x$ axis of Fig. 9d. Additionally, the $R_{wT}$ values are nearly constant at around $0.48$, which is very close to $0.52$ and can be derived from free-convective scaling relations for the vertical velocity and temperature fluctuations (Wyngaard, 2010). Therefore, by combining all these insights, one can claim that our selected periods were prototypical of a buoyancy-driven flow.
Appendix B Applicability of elliptical model
Figure 10: For a specific 30-min period (10:30-11:00 UTC, 18 May 2024) and Section-1 DTS measurements, (a) the contours of space-time correlation curves are shown where the $x$ and $y$ axes represent the temporal and spatial separations, respectively. The colorbar denotes the correlation values. (b) The space-time correlation curves are shown for two specific $r$ values in order to illustrate the computation of $\tau_{0}$. (c) The scatter plots between the separation distance $r$ and $\tau_{0}$ and $\tau_{p}$ are shown, where the straight line fits are used to compute the sweeping and convection velocities separately. (d)–(e) The scatter plots between the mean wind speed from the EC measurements and the convection and sweeping velocities are shown. The different gray colored points indicate the values from the five different horizontal sections of the DTS set-up. The dash-dotted black line indicates the 1:1 relationship. (f) The scatter plots between the turbulence intensities from EC measurements and ratios of the resultant convection speed with EC mean wind are shown for the five different horizontal sections.
As opposed to a linear model to estimate the convective speeds of turbulent structures, there also exists an elliptical model. A couple of previous studies, pertaining to atmospheric surface layer and sparse canopy flows, found the elliptical model to be more appropriate while estimating the convective speeds of turbulent structures (Han and Zhang, 2022; Everard et al., 2021). In this model, one assumes the contours of space-time correlation curves can be approximated as an ellipsoid. To test this model for a highly-heterogeneous flow, such as here, in Fig. 10a we show the contours of space-time correlation curves for the period between 10:30-11:00 UTC, corresponding to Section-1 of the DTS cable. From a visual inspection, one could say that the contours of $R_{TT}(r,\tau)$ appear to be of elliptical shape. To quantify this aspect, the elliptical model computes two velocities $U_{e}$ and $V_{e}$, which are referred to as convection and sweeping velocities, respectively.
For all practical purposes, to estimate these two quantities one needs to use two different timescales (Han and Zhang, 2022). One of those timescales is $\tau_{0}$ and are calculated from the condition, $R_{TT}(r,0)=R_{TT}(0,\tau_{0})$. The second one is $\tau_{p}$, which represents the peak positions of the space-time correlation curves at different values of $r$. The concepts of $\tau_{0}$ and $\tau_{p}$ are illustrated in Fig. 10b. The blue square represents the peak of the space-time correlation curve at a specific value of $r$, thereby denoting $\tau_{p}$. On the other hand, $\tau_{0}$ represents that timescale where $R_{TT}(r,0)$ is equal to $R_{TT}(0,\tau_{0})$. For this specific example, $r=21\ \rm m$. After defining those, one fits straight lines to quantities $\tau_{0}$ and $r$, and $\tau_{p}$ and $r$ (see Fig. 10c). If the slopes of these lines are $\alpha_{0}$ ($\tau_{0}=\alpha_{0}r$) and $\beta_{p}$ ($\tau_{p}=\beta_{p}r$), respectively, then $U_{e}$ and $V_{e}$ are computed as,
$U_{e}=\frac{\beta_{p}}{\alpha^{2}_{0}},V_{e}=\frac{\sqrt{1-\beta^{2}_{p}/\alpha^{2}_{0}}}{\alpha_{0}}.$ (19)
In Figs. 10d-e, we compare both $U_{e}$ and $V_{e}$ with the mean wind speed ($U$) from the EC measurements. This comparison is done for the along-wind cases where the gray-colored squares represent the five different sections of the DTS cable. It is clear from Fig. 10d that $U_{e}$ is weakly correlated to $U$, which is in stark contrast with Fig. 4a. The sweeping velocities, on the other hand, are consistently smaller than $U$ (Fig. 10e). In an elliptical model, the resultant convective speeds are defined as, $\sqrt{U_{e}^{2}+V_{e}^{2}}$ and if this ratio $\sqrt{U_{e}^{2}+V_{e}^{2}}/U$ is compared against the turbulent intensities of the streamwise velocity component ($\sigma_{u}/U$), a significant scatter is observed (Fig. 10f). In fact, the relationship as described by Eq. 7, cannot fully explain these variations. We, therefore, conclude that for our situation the performance of the elliptical model remains poor. It is, however, unclear whether this is a general feature of heterogeneous flows or a feature specific to our site. More experimental validation is needed to confirm this aspect, which is beyond the scope of the present work.
Appendix C Spectral conversion
Figure 11: The parametric forms of vertical velocity spectra are shown with respect to (a) frequencies (Eq. 20) and (b) wavenumbers (Eq. 22). (c)–(d) The ensemble-averaged vertical velocity spectra of Figs. 7a–b are replotted. The gray and black dash-dotted lines denote the spectral slopes $-1.25$ and $-5/3$, respectively.
We derive an analytical expression to convert the spectrum from temporal to spatial domain by using scale-dependent convective speeds. We consider the spectrum of vertical velocity, which can be expressed through the following parametric form (Kaimal and Finnigan, 1994),
$S_{ww}(f)=\frac{a}{{(1+bf)}^{5/3}}.$ (20)

### v2

@@ 3.3.4 A critical frequency limit
Figure 9: The conversion of frequencies to wavenumbers are shown for a range of $\alpha$ values. The horizontal dash-dotted line indicates the critical wavenumber corresponding to the path length of the sonic anemometer.
@@
The frequency to wavenumber conversion is a non-linear operation when $\alpha$ is non-zero (see Eq. 9). To assess the impact of $\alpha$ on frequency ($f$) to wavenumber ($\kappa$) conversion, we plot the numerical solutions of Eq. 9 at different $\alpha$ values for a specific 30-min period (Fig. 9). As shown in Fig. 9, at smaller $\alpha$ values, the $f$-$\kappa$ relationship is mostly linear, albeit, when compared to $\alpha=0$, the $\kappa$ values encompass a larger range. The non-linear effects become prominent on $f$-$\kappa$ relationship at $\alpha\geq 0.5$. Apparently, at $\alpha=0.8$, the wavenumbers estimated from Eq. 9 can be as large as $10^{6}\rm\ m^{-1}$ when converting the EC sampling frequency (10 Hz) to wavenumber (Fig. 9a). This cannot be true as it suggests that the EC systems can measure eddies comparable to the dissipation scales, which are clearly smaller than the path length (distance between the transducers) of the sonic anemometers themselves. For our sonic anemometer system, the path length is around 16 cm, and hence the wavenumbers that can be realistically resolved is around $2\pi/0.3\rm\ m^{-1}$. To avoid the aliasing effect, the minimum wavelength being resolved is assumed to be twice the path length (i.e. 30 cm). This limit is shown as a horizontal dash-dotted line in Fig. 9. If we call this minimum wavelength being set by the physical dimension of the sonic anemometer as $\kappa_{\rm cr}$, then from Eq. 9, one can convert $\kappa_{\rm cr}$ to a frequency limit ($f_{\rm cr}$) in the temporal space. Accordingly, one can estimate $f_{\rm cr}$ as,
$f_{\rm cr}=\frac{\kappa_{\rm cr}U_{C}}{2\pi{(1+\kappa_{\rm cr}z)}^{\alpha}}.$ (28)
@@
For our set up, $z=3.1$ m, $\kappa_{\rm cr}=2\pi/0.3\rm\ m^{-1}$, and $U_{C}$ can be found from the space-time correlation curves (see Fig. 4). For the along-wind cases, $U_{C}$ values are averaged over all the five sections. By replacing all these values in Eq. 28, at $\alpha=$ 0.8, $f_{\rm cr}$ appears to be $\approx$ 0.3 Hz (for both along- and perpendicular-wind periods), with some run-to-run variations since $U_{C}$ values are not fixed.
@@
If we consider the $U_{C}(\kappa z)$ values and therefore the derived wavenumbers from Eq. 9 to be correct, then it is unclear what the EC observations represent at frequencies above $f_{\rm cr}$ (i.e. $f\geq f_{\rm cr}$). This is because, beyond $f_{\rm cr}$, the eddy length scales derived from the frequencies are smaller than the physical dimension of the instrument itself. Therefore, it is difficult to associate any spatial scale with frequencies larger than $f_{\rm cr}$. From a practical standpoint, since all our results are from a single height, we cannot establish whether $z$ is the ideal scaling variable to describe $U_{c}(\kappa)$. If $z$ indeed is the correct scaling parameter, then depending on the observation height, $f_{\rm cr}$ values will change (see Eq. 28). We present our conclusions in the next section.
@@ 4 Conclusion
In this study, we have developed a comprehensive framework to test the validity of Taylor’s hypothesis and its impact on turbulence statistics over a highly heterogeneous forest clearcut surface. Particularly, our flow situation is symptomatic of a heterogeneous roughness sublayer flow, rather than a canonical atmospheric surface layer one. To execute our objectives, an extensive dataset encompassing both Distributed Temperature Sensing (DTS) and Eddy Covariance (EC) measurements were used. These datasets were collected during a five-month field campaign (May-September 2024) and were separated into two different categories. First, when the mean wind speed was parallel to the forest edge, and second, when it was perpendicular to the edge. For both of these wind scenarios, the conditions were highly turbulent with turbulence intensities of the streamwise wind fluctuations exceeding 0.5. For the parallel-wind conditions, by invoking the elliptic approximation model (he2010small; everard2021sweeping), we showed that the sweeping effects explain why the bulk convective speeds appear larger than the mean wind speed in such highly turbulent conditions. Corresponding to the perpendicular-wind scenario, we could not compute the parameters of the elliptic approximation model with sufficient statistical accuracy. This finding suggests that the accuracy of an elliptic approximation model to describe the space-time correlation curves is impacted in situations when the presence of an upstream forest canopy disturbs the flow over a clearcut location.
Nevertheless, irrespective of the wind direction, we demonstrated that the convective speeds of temperature structures at inertial subrange scales behave in a power-law fashion with increasing wavenumbers. These wavenumbers were directly computed from the DTS spatial data. Contrary to a previous study over a homogeneous atmospheric surface layer flow (cheng2017failure), the exponent of this power law was 0.8-0.8 instead of 1/3-1/3. The 1/3-1/3 exponent was theoretically derived under the assumption of homogeneous and isotropic turbulence at smaller scales of the flow. For the present case in hand, we hypothesize that there could be two reasons behind this disagreement. First, the turbulence remains anisotropic even at smaller scales of flow (chowdhuri2024quantifying). Second, as opposed to cheng2017failure, the conditions that we encounter in a forest clearcut are highly turbulent ($\sigma_{u}/U>0.5$), thereby suggesting that significant sweeping effects associated with large eddies can persist in the flow (everard2021sweeping). These sweeping effects cause modifications to the convection speeds of smaller-scale eddies and one perhaps needs larger exponents in the scale-dependent convection speeds to account for that. In fact, chowdhuri2024quantifying have shown that small-scale anisotropy depends on horizontal velocity variances (see their Fig. 6e), and therefore it is possible that its effect increases with $\sigma_{u}/U$, thereby causing deviations from the 1/3-1/3 scaling. Despite being a promising argument, it is still too early to recommend if 0.8-0.8 should replace 1/3-1/3 in highly turbulent conditions, since more field experiments and large eddy simulation studies over different surface and stability conditions are needed for further verification.
We evaluated the impact of this revised scale-dependent convection speed on the turbulence spectra of velocity and temperature, estimated from the temporal EC measurements. Instead of restricting ourselves to a specific exponent in scale-dependent convection speeds ($\alpha$), we carried out the spectral conversion on EC dataset for a range of $\alpha$ values. This was done to systematically investigate the impact of scale-dependent conversion on the EC measurements of turbulence statistics. Any non-zero $\alpha$ values yield a spectral slope of $\beta+(\beta-1)\alpha$ in the inertial subrange scales (obtained in the frequency domain), where $\beta$ is the spectral slope associated with $\alpha=0$ (i.e. no scale-dependent effects). This finding confirms the assessment of cheng2017failure that applying Taylor’s hypothesis systematically underestimates the turbulence frequency spectrum in the inertial subrange. For our flow conditions, $\beta=5/3$, and therefore when the temporal spectra are converted at an $\alpha$ value of 0.8, they show a broad region of $5/3+2/3\times 0.8\approx-1.13$ spectral slope. This spectral slope enhancement is confirmed from the spatial DTS temperature spectra. To offer an explanation behind this spectral slope enhancement at inertial subrange scales, we conduct a dimensional analysis by considering the influence of a large-eddy velocity scale, namely ${\rm TKE}^{1/2}$, on the inertial subrange statistics.
@@
For the heat flux cospectra, the same conversion yields a slope of $7/3+4/3\times 0.8\approx-1.3$. The appearance of a slope larger than 7/3-7/3 indicates an enhancement of turbulent fluxes at smaller scales of the flow. It remains a topic for future research to evaluate the impact of such enhancement on turbulent flux measurements. However, the non-linear conversion of frequencies to wavenumbers through scale-dependent convective speeds generates unrealistically large wavenumber values. The physical dimension of the sonic anemometers sets a strict constraint on the wavenumbers that can be resolved by the EC systems. In a temporal domain, this eventually results in a critical frequency limit beyond which the eddy length scales derived from the frequencies are smaller than the physical dimension of the instrument itself. This critical limit depends on the power-law exponent of the scale-dependent convective speeds. The existence of this limit dictates that there is a lack of space-time equivalence beyond this critical value.
@@
At present, these results are representative of buoyancy-driven roughness sublayer flows in a heterogeneous location and are also limited by a single observation height. One another drawback is, the estimations of the convective speeds are based on an imperfect temperature signal and we inherently assume that the same could be extended to other scalars and velocity components. It is not immediately apparent whether the structures in all these different turbulent variables move at same speeds. Therefore, for future studies, more extensive spatial data are needed to verify our framework, which ideally should span multiple heights, stability conditions (stable, neutral, and convective), surface roughness (forest canopies, grasslands, arid deserts, water-bodies, etc.), and signal types (velocity and scalars). To achieve this ambitious goal, the observations alone will not suffice and large eddy simulations would be needed to complement the DTS measurements.
@@ Acknowledgements.
In this appendix, we report the statistics of velocity and temperature fluctuations from the EC measurements corresponding to the along-wind conditions. The conclusions remain the same for perpendicular-wind conditions and hence not being shown. These statistics are used to assess the stability conditions. In practice, the Obukhov similarity parameter is commonly used to determine the atmospheric stability, but we deliberately refrained from that approach since this theory was developed for homogeneous flows and its validity is questionable over heterogeneous surfaces such as a forest clearcut. In Fig. 10a, we show the PDFs of velocity and temperature fluctuations. Typical to a buoyancy-driven flow, the PDFs of temperature fluctuations remain strongly non-Gaussian as opposed to the velocity ones ($u^{\prime}$ and $w^{\prime}$). In fact, the PDFs of temperature fluctuations can be represented through a cumulant expansion method, otherwise known as Gram-Charlier distribution (chowdhuri2019evaluation). This distribution can be expressed as:
@@
where $\mathcal{S}(T)$ is the skewness of temperature fluctuations. The non-Gaussianity in temperature statistics is also reflected in how the JPDFs between the vertical velocity and temperature fluctuations behave. From Fig. 10b, it is apparent that the JPDFs are strongly skewed towards quadrant III, which represents the downdraft motions. To further investigate the heat transport characteristics for our selected runs, we computed the flux-weighted JPDFs, expressed as, $w^{\prime}T^{\prime}P(w^{\prime},T^{\prime})$, where $P(w^{\prime},T^{\prime})$ is the JPDF between $w^{\prime}$ and $T^{\prime}$. The contours of this quantity are plotted in Fig. 10c, and the area under them is proportional to the amount of heat being transported (nakagawa1977prediction). It is immediately clear from Fig. 10c that the transport of heat is strongly dominated by the down-gradient quadrants (I and III) with very small amounts being transported in counter-gradient directions (II and IV).